OpenStudy (anonymous):

How can I find particular solution to: y" + y = (cos x)^2

OpenStudy (amistre64):

a few ways, but the "wronskian" method might be simplest

OpenStudy (amistre64):

y = yh + yp yh is the homogenous solution suc that: yh = c1y1 + c2y2 then yp can be set up as: W1 W W2 y1 0 y2 y'1 (cos x)^2 y'2 cross those

OpenStudy (anonymous):

I've used (cos x)^2 = (1 + cos2x)/2 Therefore getting: A + Bcos2x + Csin2x Let me try this way, I haven't learnt Wronskian yet

OpenStudy (amistre64):

ah, so you rewrote to use the "table"

OpenStudy (amistre64):

make sure your not using up one of your homogenous parts as a double

OpenStudy (anonymous):

sorry, what do you mean by using it as a double?

OpenStudy (amistre64):

for he solution to be linearly independant; you need to make sure you dont use the same parts more than once

OpenStudy (amistre64):

for example,if youve already got a c1 cos(2x) then you will need to introduce an extra variable into it such as; x cos(2x)

OpenStudy (anonymous):

Ahh I see. No the complimentary part is: c1cos(x) + c2sin(x). So I don't think that'll be an issue.

OpenStudy (amistre64):

good :)

OpenStudy (anonymous):

I think I got the answer. Thanks for the help :)

OpenStudy (amistre64):

yw